SLIT DISCHARGE SPECTROSCOPY OF JET-COOLED CYCLOPROPYL RADICAL IN THE ANTISYM-METRIC CH_{2} STRETCH (ν_{7}) MANIFOLD: STRUCTURES, TUNNELING BARRIER AND INTRAMOLEC-ULAR DYNAMICS

\underline{FENG DONG}, SCOTT DAVIS, and DAVID J. NESBITT, JILA, University of Colorado, Boulder, CO 80309.

High-resolution infrared spectra of jet-cooled cyclopropyl radical are reported for the first time, specifically sampling the in-phase antisymmetric CH_{2} stretch (ν_{7}) vibration. In addition to yielding first precise gas phase structural information for this radical, the spectra reveal doubling due to quantum tunneling of the lone α-CH with respect to the CCC plane, lift the degeneracy and generating lower (+) and upper (-) tunneling states. The bands clearly reveal intensity alternation (6:10 for even:odd Ka+Kc levels in the lower ← lower and 10:6 in the upper ← upper) due to H atom nuclear spin statistics, confirming that the tunneling transition state for cyclopropyl radical is of C_{2v} symmetry. However, in addition to the two predicted vibrational bands (i.e. lower ← lower and upper ← upper), a third band is observed due to IVR mixing of the upper tunneling component (ν_{7}) with a nearly isoenergetic dark state. From fractional populations in the ground and excited state, the tunneling splittings for cyclopropyl radical in the ground and excited (ν_{7}) state are estimated to be 3.2±0.3 cm^{−1} and 4.9±0.3 cm^{−1}, respectively. This indicates that stereoracemization of the α-CH radical center is a very fast process (k=2.0x10^{11} s^{−1}), and that the barrier decreases upon vibrational excitation of the CH_{2} in-phase antisymmetric stretch. The barrier height for α-CH inversion through the cyclopropyl plane is also extracted from the analysis of the tunneling dynamics with simple 1D potential energy surface.